OpenStudy (anonymous):

a hill is inclined at 18 degrees to the horizontal. If the base of the hill is at sea level find a. my height above sea level if I walk 1.2 km up the hill b. how far I have walked up the hill if I am 500 metres above sea level.

OpenStudy (anonymous):

Using simple right triangle trig, the sine is define as the ratio of the side opposite the known angle to the hypotenuse of the triangle, thus sin18=x/1.2 or x=1.2sin18=.37 km approximately (see drawing to follow)

OpenStudy (anonymous):



OpenStudy (anonymous):

For part (b) you can use the sine again since the given info involves the opposite side and the hypotenuse: sin18=500/y or y=500/sin18=1618 km approx. (again see drawing)

OpenStudy (anonymous):